Can someone please help me with the

2x – 5y = -1
x – 3y = -3

Answer by asuar010(338)

from the second equation we obtain x=-3+3y and if we plug it in in the first equation we get 2(-3+3y)-5y=-1 then -6+6y-5y=-1 so y=5 then we solve for x as x=-3+3(5)=12